# The Efficacy of Semantics-Preserving Transformations in Self-Supervised Learning for Medical Ultrasound

### v1

@@ Abstract
Data augmentation is a central component of joint embedding self-supervised learning (SSL). Approaches that work for natural images may not always be effective in medical imaging tasks. This study systematically investigated the impact of data augmentation and preprocessing strategies in SSL for lung ultrasound. Three data augmentation pipelines were assessed: (1) a baseline pipeline commonly used across imaging domains, (2) a novel semantic-preserving pipeline designed for ultrasound, and (3) a distilled set of the most effective transformations from both pipelines. Pretrained models were evaluated on multiple classification tasks: B-line detection, pleural effusion detection, and COVID-19 classification. Experiments revealed that semantics-preserving data augmentation resulted in the greatest performance for COVID-19 classification - a diagnostic task requiring global image context. Cropping-based methods yielded the greatest performance on the B-line and pleural effusion object classification tasks, which require strong local pattern recognition. Lastly, semantics-preserving ultrasound image preprocessing resulted in increased downstream performance for multiple tasks. Guidance regarding data augmentation and preprocessing strategies was synthesized for practitioners working with SSL in ultrasound.
Index Terms:
Data augmentation, Machine learning, Self-supervised learning, Ultrasound
I Introduction
@@
(a) Original images
(b) Preprocessed images
(c) BYOL pipeline
(d) AugUS-O pipeline
(e) AugUS-D pipeline
Figure 1: Examples of the preprocessing and data augmentation methods in this study. (a) Original images are from ultrasound exams. (b) Semantics-preserving preprocessing is applied to crop out areas external to the field of view. (c) The BYOL pipeline is a commonly employed data augmentation pipeline in self-supervised learning. (d) The AugUS-O pipeline was designed to preserve semantic content in ultrasound images. (e) AugUS-D is a hybrid pipeline whose construction was informed by empirical investigations into the BYOL and AugUS-O pipelines.
@@
  • Recommendations for practitioners working with unlabeled US datasets
@@
II Background
II-A Data Augmentation in Self-Supervised Learning
@@
The joint embedding class of SSL methods is characterized by the minimization of an objective function that, broadly speaking, encourages similarity of related pairs of inputs. Semantically related pairs of images (i.e., positive pairs) are sampled from unlabeled datasets according to a pairwise relationship. If the SSL pairwise relationship is satisfied for samples exhibiting the same class, SSL methods will likely improve the performance of a classifier [8]. Most joint embedding methods rely on data augmentation to define the pairwise relationship. Some studies have used meta-data or known relationships between samples to identify related pairs [9, 10, 11]; however, the availability of such information is rare. The choice of data augmentation transformations is therefore crucial, as it dictates the invariances learned [12]. However, the set of of useful invariances differs by the image modality and downstream problem(s) of interest. Despite this, studies continue to espouse the data augmentation pipeline used in the Bring Your Own Latent (BYOL) method, which includes random crops, horizontal reflection, color jitter, Gaussian blur, and solarization [13]. Random rotation is an example of a transformation not found in the BYOL pipeline that represents an important invariance relationship for many tasks in medical imaging. For example, random rotation has been applied in SSL pretraining with magnetic resonance exams of the prostate [14]. Moreover, the authors did not use BYOL’s Gaussian blur transformation because it may have rendered the images uninterpretable.
II-B Joint Embedding Self-Supervision in Ultrasound
@@
Recent studies have examined the use of joint embedding SSL methods for US interpretation tasks, such as echocardiogram view classification [15], left ventricle segmentation [16], and breast tumor classification [17]. Some have proposed positive pair sampling schemes customized for US. The USCL method and its successors explored contrastive learning methods where the positive pairs were weighted sums of images from the same US video [18, 19, 20]. Other methods have studied the use of images from the same video as positive pairs [11, 21]. In these studies, the set of transformations were a subset of the BYOL data augmentation pipeline, occasionally with different hyperparameters. Few studies have proposed US-specific data augmentation methods for SSL. A recent study by Chen et al. [22] applied BYOL and SimCLR to pretrain 3D convolutional DNNs with specialized data augmentation for lung consolidation detection in US videos, observing that temporal transformations were contributory to their problem. This study builds on previous literature by proposing and comparing domain-specific data augmentation and preprocessing method for multiple types of downstream tasks.
III Materials & Methods
III-A Datasets and Tasks
@@
TABLE I: Breakdown of the unlabeled, training, validation, and test sets in the private dataset. $x/y$ indicates the number of labeled videos in the negative and positive class for each binary classification task.
@@
The second data source is a private collection of lung ultrasound examinations, and we refer to it as LUSData. Access to this data was granted by Western University research ethics board (REB 116838) on January 28, 2021. LUSData contains videos from of parenchymal and pleural views of the lung. A subset of the parenchymal views have labels for the presence of A-lines or B-lines (i.e., the AB classification task). A-lines are reverberation artifacts that indicate normal lung tissue, while B-lines are axial artifacts that indicate fluid or thickness in the lung. A subset of the pleural views are labeled for the presence or absence of pleural effusion (i.e., the PE classification task), which is an accumulation of fluid around the lungs. A small fraction of the parenchymal views in LUSData possess bounding box labels for the pleural line artifact (i.e., the PL object detection task). Most exams in LUSData originated from a local healthcare centre, but a subset were acquired at another centre, which we adopt as an external test set. The labeled examples in the local dataset was split into training (70%), validation (15%), and test (15%) splits by patient. Table I provides the video and class counts of LUSData. Further dataset details are in Appendix A. All models in this study are trained on images, instead of on videos. Classification labels apply to every image in the video. However, individual images within each video labeled for the PL task have bounding box annotations.
III-B Semantics-Preserving Preprocessing
@@
III-C Ultrasound-specific Data Augmentation
Joint embedding SSL is effective when positive pairs contain similar information with respect to downstream tasks [8]. Several SSL studies applied to photographic or medical imaging datasets adopt the BYOL data augmentation pipeline used by Grill et al. [13]. The core aim of our study was to determine if semantics-preserving data augmentation would better equip pretrained feature extractors for downstream LUS tasks than the commonly applied BYOL pipeline.
We refer to a data augmentation pipeline as an ordered sequence of transformations, each applied with some probability. For clarity, we assign each transformation an alphanumeric identifier and express a data augmentation pipeline as an ordered sequence of identifiers. The BYOL pipeline transformations are detailed in Table II. The table also includes an estimate of the time to transform a single image. Details on how the runtime estimates were calculated are in Appendix B.
TABLE II: The sequence of transformations in the BYOL data augmentation pipeline [13]
@@ III-C Ultrasound-specific Data Augmentation
We designed the AugUS-O pipeline, which was intended to preserve semantic information in the entire US FOV while imposing nontrivial differences across invocations. The transformations in AugUS-O are listed below. Algorithmic details and parameter settings for the BYOL and AugUS-O pipelines are in Appendices C and D, respectively.
@@ 2. U01:
    Convexity Change: The shape of convex beams can vary, depending on the manufacturer, depth, and field of view of the probe. This transformation modifies the FOV shape such that the distance between $x_{1}$ and $x_{2}$ is altered, mimicking a change in $\theta_{0}$.
@@ 12. U11:
As is common in stochastic data augmentation, each transformation was applied with some probability. Table III gives the entire sequence of transformations, the probability with which each is applied. Visuals of positive pairs produced using the BYOL and AugUS-O augmentation pipelines can be found in Fig. 4a and Fig. 4b, respectively.
We conducted an informal assessment of the similarity of positive pairs. Positive pairs were produced for $50$ randomly sampled images, using both the BYOL and the AugUS-O pipelines. The pairs were presented in random order to one of the authors, who is an expert in point-of-care US. They were aware of the two pipelines, but were not told which pipeline produced each pair. The expert was asked to mark the pairs they believed conveyed the same clinical impression. We observed that $58\%$ of pairs produced with the BYOL pipeline were marked as similar, whereas $70\%$ of the AugUS-O pairs were marked as similar. While not conclusive, this check added credence to the semantics-preserving intention of the design.
III-D Discovering Semantically Contributory Transformations
A major aim of this work was to explore the utility of various data augmentation schemes during pretraining. As such, we conducted leave-one-out analysis for each of the BYOL and AugUS pipelines to estimate the impact of each transformation on models’ ability to solve downstream classification tasks. We pretrained separate models on the unlabeled portion of LUSData, using an altered version of a pipeline with one transformation omitted. We then conducted $10$-fold cross-validation on the LUSData training set for downstream classification tasks for each pretrained model. The median cross-validation test performance for each model pretrained using an ablated pipeline was compared to a baseline model that was pretrained with the entire pipeline. The experiment was conducted for both the BYOL and AugUS pipelines. Any transformations that, when omitted, resulted in worsened performance on either AB or PE were deemed contributory.
TABLE III: The sequence of transformations in the US-specific augmentation pipeline
@@ III-D Discovering Semantically Contributory Transformations
III-E Training Protocols
We adopted the MobileNetV3Small architecture [29] for all experiments in this study and pretrained using the SimCLR method [30]. MobileNetV3Small was chosen due to its real-time inference capability on mobile devices and its use in prior work by VanBerlo et al.for similar tasks [31]. Images were resized to $128\times 128$ pixels prior to the forward pass. Unless otherwise stated, backbones (i.e., feature extractors) were initialized using ImageNet-pretrained weights [32] and were pretrained using the LARS optimizer [33] with a batch size of $512$, a base learning rate of $0.2$ and a linear warmup with cosine decay schedule. Pretraining was conducted for $3$ epochs with $0.3$ warmup epochs for LUSData, and $100$ epochs with $10$ warmup epochs for COVIDx-US.
@@ III-E Training Protocols
Although this study focused on classification tasks, we also evaluated backbones on the PL object detection task using the Single Shot Detector (SSD) method [34]. SSL-pretrained backbones were used as the convolutional feature extractor. Architectural and training details for SSD are in Appendix E.
@@
IV Results
IV-A Transformation Leave-one-out Analysis
@@
A leave-one-out analysis was conducted to discover which transformations in each of the BYOL and AugUS-O pipelines were contributory to downstream task performance. We pretrained backbones using versions of each pipeline with one transformation omitted. The private LUSData training set was split by patient into $10$ disjoint subsets. For each pretrained backbone, $10$-fold cross-validation was conducted to obtain estimates of the performance of linear classifiers trained on its output feature vectors. The maximum validation area under the receiver operating curve (AUC) across epochs was recorded. Omitted transformations that resulted in statistically significant lower validation AUC for either the AB or PE task were included in a third pipeline.
@@
We conducted statistical testing to compare each of the BYOL and AugUS-O pipelines, and for each of the AB and PE tasks (described in Section III-A). Friedman’s Test for multiple comparisons [35] was conducted, with significance level $0.05$. When significant differences were found, we performed the Wilcoxon Sign-Rank Test [36] to compare the test AUCs from each ablated model to the baseline’s test AUCs. To control for false positives, the Holm-Bonferroni correction [37] was applied to keep the family-wise significance level at $0.05$.
@@
Table IV details the results of the leave-one-out analysis. Friedman’s Test detected differences in performance on both the AB and PE tasks when pretrained using the BYOL pipeline, but only the AB task exhibited differences when pretrained with the AugUS-O pipeline. As shown in Table IV, the set of transformations that exhibited statistically significant reductions in test AUC for at least one task when excluded were crop & resize (B00), color jitter (B02), CLAHE (U03), and rotation & shift (U11). Appendix F provides all test statistics from this investigation. The random crop and resize transformation (B00) Of note is the sharp decrease in performance without the random crop and resize (B00), indicating that it is a critical transformation.
@@
Using these transformations, we constructed a distilled pipeline that consists only of the above transformations. Referred hereafter to as AugUS-D, the pipeline is expressed as the following sequence: [U03, B02, U11, B00]. Fig. 4c provides some examples of positive pairs produced with AugUS-D. For more examples of pairs produced by each pipeline, see Appendix G.
(a) BYOL pipeline
(b) AugUS-O pipeline
(c) AugUS-D pipeline
@@
Figure 4: Examples of positive pairs produced using each of the BYOL, AugUS-O, and AugUS-D data augmentation pipelines.
@@
TABLE IV: A comparison of ablated versions of the BYOL and AugUS-O pipeline with one excluded transformation versus the original pipelines. Models were pretrained on the LUSData unlabeled set and evaluated on two downstream classification tasks – AB and PE. Performance is expressed as mean and median test AUC from $10$-fold cross-validation achieved by a linear classifier trained on the feature vectors of a frozen backbone.
@@
Mean Median Mean Median
BYOL None (baseline) $0.978$ $0.978$ $0.852$ $0.845$
B00 $0.864$ $0.873$ $0.695$ $0.707$
B01 $0.976$ $0.974$ $0.848$ $0.856$
B02 $0.975$ $0.975$ $0.840$ $0.842$
B03 $0.978$ $0.978$ $0.849$ $0.846$
B04 $0.976$ $0.975$ $0.840$ $0.842$
B05 $0.977$ $0.977$ $0.851$ $0.853$
AugUS-O None (baseline) $0.956$ $0.959$ $0.828$ $0.837$
U00 $0.958$ $0.957$ $0.831$ $0.839$
U01 $0.952$ $0.952$ $0.835$ $0.838$
U02 $0.965$ $0.967$§ $0.840$ $0.851$
U03 $0.950$ $0.951$ $0.825$ $0.827$
U04 $0.957$ $0.958$ $0.831$ $0.836$
U05 $0.953$ $0.952$ $0.839$ $0.845$
U06 $0.961$ $0.959$ $0.829$ $0.833$
U07 $0.959$ $0.960$ $0.838$ $0.856$
U08 $0.961$ $0.966$§ $0.834$ $0.849$
U09 $0.962$ $0.967$§ $0.838$ $0.845$
U10 $0.956$ $0.959$ $0.826$ $0.838$
U11 $0.937$ $0.939$ $0.825$ $0.823$
  • Median is significantly less than baseline.
    Median is significantly greater than baseline.
IV-B Object Classification Task Evaluation
The BYOL, AugUS-O, and AugUS-D pipelines were compared in terms of their performance on multiple downstream tasks. Model backbones were pretrained using each of the data augmentation pipelines on the union of the unlabeled and training sets in LUSData. Linear evaluation and fine-tuning experiments were performed according to the procedure explained in Section III-E. In this section, we present results on the two object classification tasks: A-line vs B-line classification (AB) and pleural effusion classification (PE).
Linear classifiers indicate the usefulness of pretrained backbones, as the only trainable weights for supervised learning are those belonging to the perceptron head. Table V reports the test set performance of linear classifiers for each task and data augmentation pipeline. On the private dataset, the AugUS-D and BYOL pipelines performed comparably well on the AB task. AugUS-D attained greater performance metrics than the BYOL pipeline on PE. To provide a visual perspective on linear classifier performance, we produced two-dimensional t-SNE embeddings of the feature vectors outputted by pretrained backbones. Shown in Fig. 5, the separability of the visual representations is consistent with linear classifier performance. For AB, AugUS-O exhibits worse separation between classes. On PE, AugUS-D appeared to have the most demarcated class separation.
TABLE V: Test set metrics for linear classification for each of the AB and PE tasks.
Task Weights Pipeline Accuracy Precision Recall AUC
AB SimCLR BYOL $\mathbf{0.932}$ $\mathbf{0.951}$ $0.819$ $0.970$

### v2

@@ Abstract
Data augmentation is a central component of joint embedding self-supervised learning (SSL). Approaches that work for natural images may not always be effective in medical imaging tasks. This study systematically investigated the impact of data augmentation and preprocessing strategies in SSL for lung ultrasound. Three data augmentation pipelines were assessed: (1) a baseline pipeline commonly used across imaging domains, (2) a novel semantic-preserving pipeline designed for ultrasound, and (3) a distilled set of the most effective transformations from both pipelines. Pretrained models were evaluated on multiple classification tasks: B-line detection, pleural effusion detection, and COVID-19 classification. Experiments revealed that semantics-preserving data augmentation resulted in the greatest performance for COVID-19 classification - a diagnostic task requiring global image context. Cropping-based methods yielded the greatest performance on the B-line and pleural effusion object classification tasks, which require strong local pattern recognition. Lastly, semantics-preserving ultrasound image preprocessing resulted in increased downstream performance for multiple tasks. Guidance regarding data augmentation and preprocessing strategies was synthesized for developers working with SSL in ultrasound.
keywords:
Data augmentation , Machine learning , Self-supervised learning , Transfer learning , Ultrasound
journal: Computers in Biology and Medicine
[…]
[uwcs]organization=Cheriton School of Computer Science, University of Waterloo, addressline=200 University Ave W, city=Waterloo, postcode=N2L 3G1, state=Ontario, country=Canada
[…]
[uwsde]organization=Department of Systems Design Engineering, University of Waterloo, addressline=200 University Ave W, city=Waterloo, postcode=N2L 3G1, state=Ontario, country=Canada
[…]
[uwomed]organization=Schulich School of Medicine and Dentistry, Western University, addressline=1151 Richmond St, city=London, postcode=N6A 3K7, state=Ontario, country=Canada
{highlights}
Novel semantics-preserving data augmentation for self-supervision in ultrasound
Semantics-preserving augmentation yielded best results for COVID-19 classification
Cropping-based data augmentation was best for object classification tasks
Semantics-preserving ultrasound preprocessing improves pretrained model performance
Guidance is provided for developers working with unlabeled ultrasound data
1 Introduction
@@
(a) (a) Original
images
(b) (b) Preprocessed images
(c) (c) StandardAug
pipeline
(d) (d) AugUS-O
pipeline
(e) (e) AugUS-D
pipeline
Figure 1: Examples of the preprocessing and data augmentation methods in this study. (a) Original images are from ultrasound exams. (b) Semantics-preserving preprocessing is applied to crop out areas external to the field of view. (c) The StandardAug pipeline is a commonly employed data augmentation pipeline in self-supervised learning. (d) The AugUS-O pipeline was designed to preserve semantic content in ultrasound images. (e) AugUS-D is a hybrid pipeline whose construction was informed by empirical investigations into the StandardAug and AugUS-O pipelines.
@@
  • Recommendations for developers working with unlabeled US datasets
@@
2 Background
2.1 Data Augmentation in Self-Supervised Learning
@@
The joint embedding class of SSL methods is characterized by the minimization of an objective function that, broadly speaking, encourages similarity of related pairs of inputs. Semantically related pairs of images (i.e., positive pairs) are sampled from unlabeled datasets according to a pairwise relationship. If the SSL pairwise relationship is satisfied for samples exhibiting the same class, SSL methods will likely improve the performance of a classifier [8]. Most joint embedding methods rely on data augmentation to define the pairwise relationship. Some studies have used meta-data or known relationships between samples to identify related pairs [9, 10, 11]; however, the availability of such information is rare. The choice of data augmentation transformations is therefore crucial, as it dictates the invariances learned [12]. However, the set of useful invariances differs by the image modality and downstream problem(s) of interest. Despite this, studies continue to espouse a data augmentation pipeline popularized by leading SSL methods, such as SimCLR [13], BYOL [14], Barlow Twins [15], and VICReg [16]. These methods utilized the same pipeline, but with minor hyperparameter variations. The pipeline includes the following transformations: random crops, horizontal reflection, colour jitter, Gaussian blur, and solarization. Hereafter, we refer to this baseline pipeline as StandardAug. Random rotation is an example of a transformation not found in the StandardAug pipeline that represents an important invariance relationship for many tasks in medical imaging. For example, random rotation has been applied in SSL pretraining with magnetic resonance exams of the prostate [17]. Moreover, the authors did not use StandardAug’s Gaussian blur transformation because it may have rendered the images uninterpretable.
2.2 Joint Embedding Self-Supervision in Ultrasound
@@
Recent studies have examined the use of joint embedding SSL methods for US interpretation tasks, such as echocardiogram view classification [18], left ventricle segmentation [19], and breast tumor classification [20]. Some have proposed positive pair sampling schemes customized for US. The USCL method and its successors explored contrastive learning methods where the positive pairs were weighted sums of images from the same US video [21, 22, 23]. Other methods have studied the use of images from the same video as positive pairs [11, 24]. In these studies, the set of transformations were a subset of the StandardAug data augmentation pipeline, occasionally with different hyperparameters. Few studies have proposed US-specific data augmentation methods for SSL. A recent study by Chen et al. [25] applied BYOL and SimCLR to pretrain 3D convolutional DNNs with specialized data augmentation for lung consolidation detection in US videos, observing that temporal transformations were contributory to their problem. This study builds on previous literature by proposing and comparing domain-specific data augmentation and preprocessing method for multiple types of downstream tasks.
3 Materials & Methods
3.1 Datasets and Tasks
@@
Table 1: Breakdown of the unlabelled, training, validation, and test sets in the private dataset. For each split, we indicate the number of distinct patients, videos, and images. $x/y$ indicates the number of labeled videos in the negative and positive class for each binary classification task. For the PL task, we indicate the number of videos with frame-level bounding box annotations.
@@
The second data source is a private collection of lung ultrasound examinations, and we refer to it as LUSData. Access to this data was granted by Western University research ethics board (REB 116838) on January 28, 2021. LUSData contains videos from of parenchymal and pleural views of the lung. A subset of the parenchymal views has labels for the presence of A-lines or B-lines (i.e., the AB classification task). A-lines are reverberation artifacts that indicate normal lung tissue, while B-lines are axial artifacts that indicate fluid or thickness in the lung. A subset of the pleural views is labeled for the presence or absence of pleural effusion (i.e., the PE classification task), which is an accumulation of fluid around the lungs. A small fraction of the parenchymal views in LUSData possess bounding box labels for the pleural line (i.e., the PL object detection task). Most exams in LUSData originated from a local healthcare centre, but a subset were acquired at another centre, which we adopt as an external test set. The labeled examples in the local dataset were split into training (70%), validation (15%), and test (15%) splits by patient. Table 1 provides the video and class counts of LUSData. Further dataset details are in A. All models in this study are trained on images, instead of on videos. Classification labels apply to every image in the video. However, individual images within each video labeled for the PL task have bounding box annotations.
3.2 Semantics-Preserving Preprocessing
@@
3.3 Ultrasound-specific Data Augmentation
Joint embedding SSL is effective when positive pairs contain similar information with respect to downstream tasks [8]. Several SSL studies applied to photographic or medical imaging datasets adopted variations of the StandardAug data augmentation pipeline. The core aim of our study was to determine if semantics-preserving data augmentation would better equip pretrained feature extractors for downstream LUS tasks than the commonly applied StandardAug pipeline.
We refer to a data augmentation pipeline as an ordered sequence of transformations, each applied with some probability. For clarity, we assign each transformation an alphanumeric identifier and express a data augmentation pipeline as an ordered sequence of identifiers. The StandardAug pipeline transformations are detailed in Table 2. The table also includes an estimate of the time to transform a single image. Details on how the runtime estimates were calculated are in B.
Table 2: The sequence of transformations in the StandardAug data augmentation pipeline
@@ 3.3 Ultrasound-specific Data Augmentation
We designed the AugUS-O pipeline, which was intended to preserve semantic information in the entire US FOV while imposing nontrivial differences across invocations. The transformations in AugUS-O are listed below.
@@ 2. U01:
    Convexity Change: The shape of convex FOVs can vary, depending on the manufacturer, depth, and field of view of the probe. This transformation modifies the FOV shape such that the distance between $x_{1}$ and $x_{2}$ is altered, mimicking a change in $\theta_{0}$.
@@ 12. U11:
Refer to Figure 3 for a visual example of each transformation in AugUS-O. Algorithmic details and parameter settings for the StandardAug and AugUS-O pipelines are in Appendices C and D, respectively. As is common in stochastic data augmentation, each transformation was applied with some probability. Table 3 gives the entire sequence of transformations, the probability with which each is applied. Visuals of positive pairs produced using the StandardAug and AugUS-O augmentation pipelines can be found in Fig. LABEL:subfig:aug-examples-1 and Fig. LABEL:subfig:aug-examples-2, respectively.
Table 3: The sequence of transformations in the US-specific augmentation pipeline.
@@
We conducted an informal assessment of the similarity of positive pairs. Positive pairs were produced for $50$ randomly sampled images, using both the StandardAug and the AugUS-O pipelines. The pairs were presented in random order to one of the authors, who is an expert in point-of-care US. They were aware of the two pipelines but were not told which pipeline produced each pair. The expert was asked to mark the pairs they believed conveyed the same clinical impression. We observed that $58\%$ of pairs produced with the StandardAug pipeline were marked as similar, whereas $70\%$ of the AugUS-O pairs were marked as similar. While not conclusive, this manual evaluation added credence to the semantics-preserving intention of the design.
3.4 Discovering Semantically Contributory Transformations
A major aim of this work was to explore the utility of various data augmentation schemes during pretraining. As such, we conducted leave-one-out analysis for each of the StandardAug and AugUS pipelines to estimate the impact of each transformation on models’ ability to solve downstream classification tasks. We pretrained separate models on the unlabeled portion of LUSData, using an altered version of a pipeline with one transformation omitted. We then conducted $10$-fold cross-validation on the LUSData training set for downstream classification tasks for each pretrained model. The median cross-validation test performance for each model pretrained using an ablated pipeline was compared to a baseline model that was pretrained with the entire pipeline. The experiment was conducted for both the StandardAug and AugUS pipelines. Any transformations that, when omitted, resulted in worsened performance on either AB or PE were deemed contributory.
3.5 Training Protocols
We adopted the MobileNetV3Small architecture [32] for all experiments in this study and pretrained using the SimCLR method [13]. MobileNetV3Small was chosen due to its real-time inference capability on mobile devices and its use in prior work by VanBerlo et al.  for similar tasks [33]. The SimCLR projector was a $2$-layer multilayer perceptron with $576$ nodes per layer. Images were resized to $128\times 128$ pixels prior to the forward pass. Unless otherwise stated, backbones (i.e., feature extractors) were initialized using ImageNet-pretrained weights [34] and were pretrained using the LARS optimizer [35] with a batch size of $1024$, a base learning rate of $0.2$ and a linear warmup with cosine decay schedule. Pretraining was conducted for $3$ epochs with $0.3$ warmup epochs for LUSData, and $100$ epochs with $10$ warmup epochs for COVIDx-US.
@@ 3.5 Training Protocols
Although this study focused on classification tasks, we also evaluated backbones on the PL object detection task using the Single Shot Detector (SSD) method [36]. SSL-pretrained backbones were used as the convolutional feature extractor. Architectural and training details for SSD are in E.
@@
4 Results
4.1 Transformation Leave-one-out Analysis
@@
A leave-one-out analysis was conducted to discover which transformations in each of the StandardAug and AugUS-O pipelines were contributory to downstream task performance. We pretrained backbones using versions of each pipeline with one transformation omitted. The private LUSData training set was split by patient into $10$ disjoint subsets. For each pretrained backbone, $10$-fold cross-validation was conducted to obtain estimates of the performance of linear classifiers trained on its output feature vectors. The maximum validation area under the receiver operating curve (AUC) across epochs was recorded. Omitted transformations that resulted in statistically significant lower validation AUC for either the AB or PE task were included in a third pipeline.
@@
We conducted statistical testing to compare each of the StandardAug and AugUS-O pipelines, and for each of the AB and PE tasks (described in Section 3.1). Friedman’s Test for multiple comparisons [37] was conducted, with significance level $0.05$. When significant differences were found, we performed the Wilcoxon Sign-Rank Test [38] to compare the test AUCs from each ablated model to the baseline’s test AUCs. To control for false positives, the Holm-Bonferroni correction [39] was applied to keep the family-wise significance level at $0.05$.
@@
Table 4 details the results of the leave-one-out analysis. Friedman’s Test detected differences in performance on both the AB and PE tasks when pretrained using the StandardAug pipeline, but only the AB task exhibited differences when pretrained with the AugUS-O pipeline. As shown in Table 4, the set of transformations that exhibited statistically significant reductions in test AUC for at least one task when excluded were crop & resize (B00), color jitter (B02), CLAHE (U03), and rotation & shift (U11). F provides all test statistics from this investigation. The random crop and resize transformation (B00) Of note is the sharp decrease in performance without the random crop and resize (B00), indicating that it is a critical transformation.
@@
Using these transformations, we constructed a distilled pipeline that consists only of the above transformations. Referred hereafter to as AugUS-D, the pipeline is expressed as the following sequence: [U03, B02, U11, B00]. Fig. LABEL:subfig:aug-examples-3 provides some examples of positive pairs produced with AugUS-D. For more examples of pairs produced by each pipeline, see G.
@@
Figure 4: Examples of positive pairs produced using each of the StandardAug, AugUS-O, and AugUS-D data augmentation pipelines.
@@
Table 4: A comparison of ablated versions of the StandardAug and AugUS-O pipeline with one excluded transformation versus the original pipelines. Models were pretrained on the LUSData unlabeled set and evaluated on two downstream classification tasks – AB and PE. Performance is expressed as mean and median test AUC from $10$-fold cross-validation achieved by a linear classifier trained on the feature vectors of a frozen backbone.
@@
Mean (std) Median Mean Median
StandardAug None $0.978\,(0.007)$ $0.978$ $0.852\,(0.040)$ $0.845$
B00 $0.864\,(0.022)$ $0.873$ $0.695\,(0.050)$ $0.707$
B01 $0.976\,(0.006)$ $0.974$ $0.848\,(0.046)$ $0.856$
B02 $0.975\,(0.007)$ $0.975$ $0.840\,(0.046)$ $0.842$
B03 $0.978\,(0.007)$ $0.978$ $0.849\,(0.044)$ $0.846$
B04 $0.976\,(0.007)$ $0.975$ $0.840\,(0.041)$ $0.842$
B05 $0.977\,(0.007)$ $0.977$ $0.851\,(0.041)$ $0.853$
AugUS-O None $0.956\,(0.013)$ $0.959$ $0.828\,(0.030)$ $0.837$
U00 $0.958\,(0.011)$ $0.957$ $0.831\,(0.034)$ $0.839$
U01 $0.952\,(0.016)$ $0.952$ $0.835\,(0.027)$ $0.838$
U02 $0.965\,(0.011)$ $0.967$§ $0.840\,(0.032)$ $0.851$
U03 $0.950\,(0.011)$ $0.951$ $0.825\,(0.028)$ $0.827$
U04 $0.957\,(0.013)$ $0.958$ $0.831\,(0.034)$ $0.836$
U05 $0.953\,(0.014)$ $0.952$ $0.839\,(0.024)$ $0.845$
U06 $0.961\,(0.009)$ $0.959$ $0.829\,(0.037)$ $0.833$
U07 $0.959\,(0.012)$ $0.960$ $0.838\,(0.035)$ $0.856$
U08 $0.961\,(0.013)$ $0.966$§ $0.834\,(0.027)$ $0.849$
U09 $0.962\,(0.012)$ $0.967$§ $0.838\,(0.030)$ $0.845$
U10 $0.956\,(0.011)$ $0.959$ $0.826\,(0.035)$ $0.838$
U11 $0.937\,(0.020)$ $0.939$ $0.825\,(0.028)$ $0.823$
  • Median is significantly less than baseline, where no transformations were omitted.
    Median is significantly greater than baseline, where no transformations were omitted.
4.2 Object Classification Task Evaluation
The StandardAug, AugUS-O, and AugUS-D pipelines were compared in terms of their performance on multiple downstream tasks. Model backbones were pretrained using each of the data augmentation pipelines on the union of the unlabeled and training sets in LUSData. Linear evaluation and fine-tuning experiments were performed according to the procedure explained in Section 3.5. In this section, we present results on the two object classification tasks: A-line vs B-line classification (AB) and pleural effusion classification (PE).
Linear classifiers indicate the usefulness of pretrained backbones, as the only trainable weights for supervised learning are those belonging to the perceptron head. Table 5 reports the test set performance of linear classifiers for each task and data augmentation pipeline. On the private dataset, the AugUS-D and StandardAug pipelines performed comparably well on the AB task. AugUS-D attained greater performance metrics than the StandardAug pipeline on PE. To provide a visual perspective on linear classifier performance, we produced two-dimensional t-SNE embeddings of the feature vectors outputted by pretrained backbones. Shown in Fig. 5, the separability of the visual representations is consistent with linear classifier performance.
Table 5: Test set performance for linear classification (LC) and fine-tuning (FT) experiments with the AB and PE tasks. Binary metrics are averages across classes. The best observed metrics in each experimental setting are in boldface.
Train Setting Task Initial Weights Pipeline Accuracy Precision Recall AUC
LC AB SimCLR StandardAug $\mathbf{0.932}$ $\mathbf{0.951}$ $0.819$ $0.970$